# Quantum-Classical Hybrid Quantized Neural Network

### v1

@@ 4 QCBO Model of Neural Network Training
where under sample $x$, $y_{x}$ represents the label, $a^{L}_{x}$ denotes the output of the last layer; $a^{l}_{x,j}$ denotes the output of the $j$-th neuron of the $l$-th layer; $C(M_{i},y_{x})$ indicates the the loss when the last layer’s output is in the $i$-th subinterval and the target is $y_{x}$. For any variable $\tau$, $\beta_{\tau}=1$ if $\tau$ lies within the $i$-th sub-interval and 00 otherwise.
Note that $w^{l}_{jk}a^{l-1}_{x,k}$ contains quadratic terms, which will introduce higher-order terms when transformed into QUBO form. Therefore, it is necessary to perform a reduction operation to lower the order. For example, when using the penalty method to obtain QUBO model, we can replace the term $x_{1}x_{2}$ by binary variable $y$ and add the Rosenberg polynomial [39] $h(x_{1},x_{2},y)=3y+x_{1}x_{2}-2y(x_{1}+x_{2}),$ to the objective function, then the order can be reduced. Rosenberg polynomial is a method for simplifying higher-order terms in a QUBO expression by introducing auxiliary variables. These auxiliary variables help to linearize the expression, making it easier to handle.
In addition, the non-binary variables $a^{l}_{x,j}$, $w^{l}_{jk}$, $b^{l}_{j}$ are represented using binary encoding as following: $a^{l}_{x,j}=\sum_{\ell}2^{\ell}\cdot\delta^{(\ell)}_{a^{l}_{x,k}},w^{l}_{jk}=\sum_{\ell}2^{\ell}\cdot\delta^{(\ell)}_{w^{l}_{jk}},b^{l}_{j}=\sum_{\ell}2^{\ell}\cdot\delta^{(\ell)}_{b^{l}_{j}}.$
In Appendix, we also provide the QCBO model when spline approximation with piecewise linear functions are used.
Model Complexity.
The table below summarizes the required number of ising spins for various variables in our quadratic programming model, showing that the overall number of ising spins required is $O(nWDN)$, for handling datasets with $N$ samples and neural network with width $W$ and depth $D$.
@@
Table 1: Summary of Required Number of ising spins for Each Variable
Note that the number of samples $N$ can potentially be very large. However, it can be bounded while ensuring the efficiency of our learning process, this is where the concept of VC dimension becomes crucial.
@@
Note that in the PAC (Probably Approximately Correct) learning framework, the number of samples can be bounded as
$N=O\Big{(}\frac{\text{VCdim}}{\varepsilon}\log\frac{1}{\delta}\Big{)}=O\Big{(}\frac{W^{2}D(\log W+\log D)}{\varepsilon}\log\frac{1}{\delta}\Big{)}$ (20)
5 CP Framework and Hybrid algorithm
@@
In this section, we confront the formidable optimization challenges stemming from the proliferation of constraints in large scale network in the aforementioned model. The abundance of constraints not only complicates the optimization process but also necessitates the determination of numerous penalty coefficients in traditional penalty method. The task of figuring out these penalty coefficients becomes increasingly intricate as the number of constraints grows, leading to potential inefficiencies and suboptimal solutions.
To surmount this hurdle, we use the Quantum Conditional Gradient Descent (QCGD) algorithm developed in [42], a cutting-edge hybrid classical-quantum framework tailored to address quadratic, linearly-constrained, binary optimization problems on quantum machines. Before introducing QCGD, we first introduce its theoretical foundation, the copositive programming framework.
5.1 Copositive Programming Framework and Order Reduction
Copositive Programming (CP) [43] is a mathematical optimization branch that has garnered significant attention for its applications in various optimizartion problems, particularly in the context of combinatorial optimization. In copositive programming, the objective is to optimize over the cone of copositive matrices, which are matrices for which the quadratic form yields nonnegative values over the nonnegative orthant. One key application of copositive programming is in transforming binary quadratic problems into completely positive problems [43]. This transformation involves reformulating the standard quadratic problem as a CP program, enabling the optimization of a quadratic form over the Cartesian product of simplices.
$\min\mathbf{x}^{T}\mathbf{Q}\mathbf{x}$ (21)
$s.t.\mathbf{A}\mathbf{x}=\mathbf{b}.$ (22)
According to [43, 42], it can be converted to the a completely positive program.
Order reduction.
$\min\;$ $\mathbf{x}^{T}\mathbf{Q}_{0}\mathbf{x}$ (23)
$s.t.\;$ $\mathbf{A}\mathbf{x}=\mathbf{b}$ (24)
$\mathbf{x}^{T}\mathbf{Q}_{t}\mathbf{x}\leq c_{t},1\leq t\leq T$ (25)
Improved Linearization Procedure.
Consider the term $z^{l}_{x,j,k}=w^{l}_{jk}a^{l-1}_{x,k}$ in QCBO model, which can be expressed as $z^{l}_{x,j,k}=\sum_{\ell=1}^{\bar{\ell}}2^{\ell}\cdot w^{l}_{jk}\cdot\delta^{(\ell)}_{a^{l}_{x,k}}$. In the context of neural network training, quadratic constraints arise from the $w^{l}_{jk}a^{l-1}_{x,k}$ terms. Note that the bilinear components involving $w^{l}_{jk}$ and $a^{l-1}_{x,k}$ share the common variable $w^{l}_{jk}$. To linearize these terms effectively, we can use the following constraints:
$z^{l}_{x,j,k}\leq(2^{\bar{\ell}+1}-2)\cdot w^{l}_{jk}.$ (28)
Note that here we treat $w^{\ell}_{jk}$ as binary variables, for integer-valued $w^{\ell}_{jk}$, the linearization process described above is still applicable. We only need to perform the same operation on the bits of the binary encoding of $w^{\ell}_{jk}$ and $a^{l-1}_{x,k}$.
5.2 Quantum Conditional Gradient Descent Algorithm With Inexact Oracle
@@
Our workflow starts with data being fed into the neural network. The optimization problem is then formulated as a QCBO. This problem is subsequently linearized and transformed into a copositive programming problem. In the hybrid quantum-classical iterative process, during each iteration, quantum conditional gradient method calculates the update direction by minimizing a linear approximation of the penalized proxy of the objective function. QCGD formulates the linear minimization problem as a QUBO sub-problem, which is suitable for Ising machines. CIM solves the QUBO problem and the solution obtained is fed back into the classical computing step to calculate the direction of iteration. This iterative loop continues, with the QUBO matrix being updated at each step, until convergence towards an optimal solution is achieved.
@@
  2. 2.
  3. 3.
    Form the gradient matrix $\mathbf{G}_{t}$ using the coefficient matrix in copositive programming formulation and constraint errors above. To ensure convergence, solve the subproblem $\min_{\mathbf{w}\in\mathbb{Z}_{2}^{p}}\mathbf{w}^{\top}\mathbf{G}_{t}\mathbf{w}$ $m_{t}$ times on a quantum annealer, selecting the best solution as:
  4. 4.
  5. 5.
  6. 6.
Ising machine as an (inexact) oracle.
The practical implementation of QCGD algorithm is challenged by the error rates associated with quantum computers, leading to inaccuracies in the computation results. This inaccuracy is particularly pronounced in the oracle, which plays a crucial role in the QCGD algorithm by providing direction in each iteration. The consideration of the inexact oracle in the context of QCGD is therefore crucial for several reasons. Firstly, due to the inherent limitations of quantum hardware, achieving perfect precision in quantum computations is challenging. The inexactness in the oracle reflects these limitations and underscores the need to develop algorithms that are robust to such imperfections. Secondly, the use of an inexact oracle also introduces a trade-off between computational accuracy and efficiency.
The key insight of copositive programming framework is the quadratic objective and linear constraints of quadratic binary optimization problem can be represented as linear functions of a copositive matrix.
Theorem 3
Consider the Quadratic Constrained Binary Optimization (QCBO) problem solved using the QCGD algorithm with a random quantum oracle. At step $t$, the error in the objective in a single run of the quantum oracle, has mean $\mu_{t}$ and variance $\sigma_{t}$, both of which have time-independent upper bounds. Then:
  1. 1.
    QCGD algorithm converges to the optimal solution almost surely, i.e.,
@@
    The total time to solution satisfies:
    $\text{TTS}_{\text{QCBO}}=O\left(\text{TTS}_{\text{QUBO}}\cdot T\cdot\log n\right),$ (32)
@@
  • $\text{TTS}_{\text{QUBO}}=R\tau$ be the time to solution for a single QUBO subproblem, where $\tau$ is the time per quantum oracle call, $R=\left\lceil\frac{\log(1-p_{R})}{\log(1/(1-p_{\tau}))}\right\rceil$ is the number of samples needed to achieve probability $p_{R}$ of finding the optimum.
Before presenting the proof, we first introduce the definition of inexact oracle in each step of QCGD.
Definition 1 ($(\delta,\varepsilon)$-Inexact oracle [44, 45])
For problem
$\min_{\mathbf{D}}\mathrm{Tr}(\mathbf{Q}^{(t)}_{\text{QUBO}}\mathbf{D})$ (33)
A quantum computer without any imprecisions corresponds to the scenario where $\delta=1$ and $\varepsilon=0$. The following lemma indicates that convergence can still be achieved even if the single quantum computation acts as an inexact oracle. This theoretical framework ensures that the overall algorithm remains robust despite potential inaccuracies in the quantum computation. Theoretical proof follows from that in [42], and we include it in Appendix for completeness.
Proposition 1
[…]
Theorem 3 The subproblem in the $t$-th iteration is solved via $m_{t}=c\log n$ repetitions, selecting the best solution among samples $\{\mathbf{D}^{(i)}_{t}\}_{1\leq i\leq m_{t}}$, i.e., $\mathbf{D}_{t}=\mathbf{D}^{(i_{t})}_{t}$, where
@@
which is independent of iteration index $t$, assuming each element in $\mathbf{Q}^{(t)}_{\text{QUBO}}$ is constant. According to Proposition 1, we can get the order of convergence in expectation, and the almost sure convergence follows from the concentration probability bound that converges to $1$ when $T$ is infinty.
The bound on TTS can be also derived from the concentration result in Proposition 1. A more straightforward approach is, based on the union bound, the probability that
@@
Therefore, With high probability $1-\frac{1}{\mathrm{poly}(n)}\geq p_{R}$, $\mathbf{D}_{t}=\mathbf{D}^{*}_{t}$ for all $t$, yielding the convergence to optimal solution. As per-iteration time of QCGD is $m_{t}\tau=O(\log n\cdot\text{TTS}_{\text{QUBO}})$, we can obtain the second conclusion for $\text{TTS}_{\text{QCBO}}$.
Convergence with limited precision.
Quantum computers, such as those implementing adiabatic quantum computing (AQC) for solving QUBO problems, face challenges due to limited parameter precision. In theory, QUBO parameters are real-valued, but practical hardware—whether quantum annealers like D-Wave’s or other accelerators—relies on finite precision representations, such as fixed-bit integers [46]. This limitation introduces perturbations in quantum systems, which distort the energy landscape and can shift the global optimum.
@@
Let the coefficients of the QUBO problem be truncated to $d=O(\log n)$ digits. If QCGD is used with these truncated coefficients, it will retains convergence guarantees and the number of iterations remains in the same order.
@@
The formal proof is presented in Appendix. This result is appealing because it demonstrates that even with limited precision, the hybrid algorithm can still achieve convergence without a substantial increase in computational effort. This behavior can be interpreted as a form of implicit error correction. The introduced errors are effectively controlled and do not accumulate in a way that disrupts the convergence properties of the algorithm.
Lazy implementation of Quantum Conditional Gradient Descent Algorithm.
We can determine whether solving the QUBO problem is necessary based on the spectral gap:
then it implies that the perturbations in the solution landscape are insignificant. Consequently, the algorithm can bypass solving the latest QUBO problem. By leveraging this stability, potentially through a threshold-based criterion to detect negligible changes, computational overhead can be reduced, particularly in resource-constrained environments.
6 Experimental Results
@@
In this section we present an experimental study on the effectiveness of using CIM to train a neural network for classification of coat and sandal images from the Fashion MNIST dataset [47]. The dataset consists of $60,000$ training images and $10,000$ test images, each representing a grayscale clothing item from $10$ classes, including T-shirts/tops, trousers, pullovers, dresses, coats, sandals, shirts, sneakers, bags, and ankle boots. Each image is $28\times 28$ pixels, with pixel values ranging from 00 to $255$. Additional details of the experimental results are deferred to Appendix.
6.1 Coherent Ising Machine: Performance and Classical Benchmarks
As shown in Figure 4, the coherent Ising machine (CIM) operates by leveraging optical parametric oscillators (OPOs) to solve combinatorial optimization problems. The system begins with a continuous wave (CW) laser at 1560 nm, which is modulated and amplified to produce pump pulses. These pulses undergo phase-sensitive amplification in a loop containing key components such as intensity modulators, a delay line, a second harmonic generation (SHG) unit, and a periodically poled lithium niobate (PPLN) crystal. The generated signal and idler photons are then processed using balanced homodyne detection (BHD), with feedback control provided by an FPGA module. This setup enables the CIM to map optimization problems onto the energy landscape of the Ising model, efficiently finding optimal or near-optimal solutions through the coherent properties of the optical signals.
@@ 6.1 Coherent Ising Machine: Performance and Classical Benchmarks
Figure 5 illustrates the energy evolution during the optimization process using CIM to directly solve the neural network training problem. CIM identifies the optimal solution at the 88th iteration, reaching the optimal Hamiltonian value in 0.185 ms, with the resulting model achieving a test accuracy of $94.85\%$.
Figure 5: Energy evolution curve during the optimization of a quantized neural network using a coherent Ising machine (CIM).
@@
To evaluate the performance of our quantum-based approach for training quantized neural networks, we present a series of comparative analyses. Table 2 compares the test accuracy and running time of our model and quantum approach against established classical approaches, including Straight Through Estimator (STE) [10, 6] and BinaryConnect [48]. We provide a brief introduction of the two classical algorithms in Appendix G. The proposed quantum-based method achieves a test accuracy of $94.95\%$ with a significantly reduced running time of $0.185$ ms, demonstrating superior efficiency and effectiveness.
[…]
STE [10, 6] 93.90% 6.1215 s
BinaryConnect [48] 94.74% 6.5204 s
This paper 94.95% 0.185 ms
Table 2: Comparison of Test Accuracy and Running Time
@@
Further, Table 3 examines the memory demand and inference complexity, contrasting a neural network with continuous weights and sigmoid activation functions against our quantized approach, while approximating the activation function with a simple piecewise linear function. Our method reduces memory usage to 2.75 bytes and inference time to 0.0762 seconds while maintaining a test accuracy of $94.95\%$, highlighting its resource efficiency.
[…]
Continuous Weights & Sigmoid 44 0.2115 97.10
This Paper 2.75 0.0762 94.95
Table 3: Comparison of Memory and Inference Complexity
These results collectively underscore the advantages of leveraging quantum computing for training quantized neural networks, offering improvements in speed, and resource efficiency.
Robustness of spline interpolation.
For spline interpolation, the number of intervals impacts qubit requirements and must align with quantum hardware capacity. In our experiments, we divided the range into four intervals to balance qubit use and segmentation fidelity. Results for breakpoints $[-8,-c,0,c,8]$ are summarized below:
@@ 6.2 Convergence Results of the Hybrid Algorithm
Figure 6 presents the convergence behavior of the Quantum Conditional Gradient Descent algorithm applied to the neural network training problem, utilizing a hybrid quantum-classical approach with CIM and Gurobi solvers. The algorithm converges to the optimal solution in 321 iterations, with the objective residual and constraint residual plotted on a log-log scale. The CIM solver demonstrates a steady decline, reducing the objective and constraint residual below a small value, while Gurobi continues to converge but with oscillations in the constraint residual between iterations 30 and 100. Notably, the hybrid strategy achieves convergence in just 321 iterations, demonstrating its computational efficiency, by effectively combining the strengths of CIM and classical approach. To further validate the robustness and reproducibility of these results, five repeated experiments were conducted, achieving convergence to the optimal solution in 723, 338, 263, 432, and 427 iterations, respectively. These consistent outcomes across multiple trials highlight the algorithm’s reliable performance and repeatability.
6.3 Robust of Hybrid Approach: Impact of Coefficient Matrix Rounding
Convergence under rounded matrix.
Figure 7: Circular heatmaps depict the evolution of the QUBO matrix during the QCGD algorithm for different significant digit precisions (1 to 6 digits) across selected iterations (300, 600, 900, 1200, 1500, 1800, 2100, 2400, 2700, and 3000). The color intensity represents matrix element values, with the color bar indicating the range from 0.0 (purple) to 1.0 (yellow). Each row corresponds to a specific precision level, illustrating the iterative trajectory and convergence behavior.
To investigate the convergence behavior of the hybrid algorithm under varying levels of matrix rounding for the neural network training problem, the matrix coefficients were systematically rounded to digit precisions ranging from 1 to 10, and the algorithm was executed until convergence to optimal. The matrix elements were rounded to 1 through 10 digits, and the number of iterations required to reach the optimal solution for each precision level was as follows:
@@ 6.3 Robust of Hybrid Approach: Impact of Coefficient Matrix Rounding
These results indicate variability in convergence speed, with lower precision levels (e.g., 1-3 digits) requiring substantially more iterations (up to 43,343 for 3 digits), likely due to increased numerical instability and larger initial differences. In contrast, higher precision levels (e.g., 5-10 digits) converge more efficiently, with the minimum iterations observed at 2904 for 10 digits, suggesting that increased precision enhances convergence stability.

### v2

@@ Results
where $x$, $y_{x}$ represents the label, $a^{L}_{x}$ denotes the output of the last layer; $a^{l}_{x,j}$ denotes the output of the $j$-th neuron of the $l$-th layer; $C(M_{i},y_{x})$ indicates the the loss when the last layer’s output is in the $i$-th subinterval and the target is $y_{x}$. For any variable $\tau$, $\beta_{\tau}=1$ if $\tau$ lies within the $i$-th sub-interval and 00 otherwise.
Note that $w^{l}_{jk}a^{l-1}_{x,k}$ contains quadratic terms, which will introduce higher-order terms when transformed into the QUBO form. Therefore, it is necessary to perform a reduction operation to the lower order. For example, when using the penalty method to obtain the QUBO model, we can replace the term $x_{1}x_{2}$ by a binary variable $y$ and add the Rosenberg polynomial [41] $h(x_{1},x_{2},y)=3y+x_{1}x_{2}-2y(x_{1}+x_{2}),$ to the objective function. These auxiliary variables help to linearize the expression, making it easier to handle.
In addition, the non-binary variables $a^{l}_{x,j}$, $w^{l}_{jk}$, $b^{l}_{j}$ are represented using binary encoding as following: $a^{l}_{x,j}=\sum_{\ell}2^{\ell}\cdot\delta^{(\ell)}_{a^{l}_{x,k}},w^{l}_{jk}=\sum_{\ell}2^{\ell}\cdot\delta^{(\ell)}_{w^{l}_{jk}},b^{l}_{j}=\sum_{\ell}2^{\ell}\cdot\delta^{(\ell)}_{b^{l}_{j}}.$ In supplementary, we also provide the details of the QCBO model when spline approximation with piecewise linear functions is used.
Spin Complexity and Upper Bounds.
Table 1 summarizes the required number of ising spins for various variables in our quadratic programming model, showing that the overall number of ising spins required is $O(nWDN)$, for handling the datasets with $N$ samples and the neural network with width $W$ and depth $D$.
@@
Table 1: The number of required ising spins for each variable
Recall that our objective is to minimize:
$\mathcal{L}(\theta)=\mathbb{E}_{(x,y_{x})\sim\mathcal{D}}[C_{x}(f_{\theta}(x),y_{x})],$ (12)
where $\theta=\{w^{l},b^{l}\}_{l=1}^{L}$, $f_{\theta}(x)=a^{L}_{x}$. We solve empirical risk minimization (ERM) over $N$ i.i.d. samples $\{(x_{i},y_{i})\}_{i=1}^{N}\sim\mathcal{D}$:
$\hat{\theta}_{N}=\arg\min_{\theta\in\Theta}\hat{\mathcal{L}}_{N}(\theta)=\frac{1}{N}\sum_{i=1}^{N}C_{x_{i}}(f_{\theta}(x_{i}),y_{i}).$
Note that the number of samples $N$ can be bounded while ensuring the efficiency of our learning process, this is where the concept of VC dimension becomes crucial.
@@
Within the Probably Approximately Correct (PAC) learning framework [44], the focus is on controlling the generalization error [44], a measure of how accurately an algorithm is able to predict outcomes for previously unseen data. This framework ensures that, with high probability $1-\alpha$, the generalization error is confined within a precision $\varepsilon$. Consequently, the required number of samples can be bounded as:
$N=O\Big{(}\frac{\text{VCdim}}{\varepsilon}\log\frac{1}{\alpha}\Big{)}=O\Big{(}\frac{W^{2}D(\log W+\log D)}{\varepsilon}\log\frac{1}{\alpha}\Big{)}.$ (13)
Moreover, from an optimization perspective, we can derive a sample complexity bound tailored to the optimization error, defined as the discrepancy between the expected loss of the learned parameters and the minimum loss, i.e., $\mathcal{L}(\hat{\theta}_{N})-\mathcal{L}(\theta^{*})$. This focus on optimization error is particularly pertinent given that our study is about addressing optimization problems,
Theorem 2
For a quantized neural network with $B$-bit parameters, let $\varepsilon,\alpha\in(0,1)$, and assume the loss function $C_{x}(\cdot)$ is bounded in $[0,C_{\max}]$. Then, with probability at least $1-\alpha$,
$\mathcal{L}(\hat{\theta}_{N})\leq\mathcal{L}(\theta^{*})+\varepsilon,$ (14)
provided $N=O\left(\frac{C_{\max}^{2}}{\varepsilon^{2}}\log\frac{|\Theta|}{\alpha}\right)$. Substituting $|\Theta|$:
$N=O\left(\frac{C_{\max}^{2}}{\varepsilon^{2}}\sum_{l=1}^{L}d_{l}(d_{l-1}+1)B\log 2+\log\frac{1}{\alpha}\right),$ (15)
where $d_{l}$ and $d_{l-1}$ are the layer dimensions.
Proof: Define the function class $\mathcal{G}=\{g_{\theta}(x,y)=C_{x}(f_{\theta}(x),y):\theta\in\Theta\}$. Given the finite $\Theta$, apply Hoeffding’s inequality for $g_{\theta}\in[0,C_{\max}]$:
$P\left(\left|\frac{1}{N}\sum_{i=1}^{N}g_{\theta}(x_{i},y_{i})-\mathbb{E}_{(x,y_{i})}[g_{\theta}(x_{i},y_{i})]\right|>\varepsilon/2\right)\leq 2\exp\left(-\frac{2N\varepsilon^{2}}{C_{\max}^{2}}\right).$ (16)
Using a union bound over all $|\Theta|$ functions:
$P\left(\sup_{\theta\in\Theta}\left|\frac{1}{N}\sum_{i=1}^{N}g_{\theta}(x_{i},y_{i})-\mathbb{E}_{(x,y_{i})}[g_{\theta}(x_{i},y_{i})]\right|>\varepsilon/2\right)\leq 2|\Theta|\exp\left(-\frac{2N\varepsilon^{2}}{C_{\max}^{2}}\right).$ (17)
Set the probability to $\alpha$, i.e., $2|\Theta|\exp\left(-\frac{2N\varepsilon^{2}}{C_{\max}^{2}}\right)\leq\alpha$, we have $N\geq\frac{C_{\max}^{2}}{2\varepsilon^{2}}\log\frac{2|\Theta|}{\alpha}$. Since $\hat{\theta}_{N}$ minimizes ^$\hat{\mathcal{L}}_{N}$:
$\mathcal{L}(\hat{\theta}_{N})\leq\hat{\mathcal{L}}_{N}(\hat{\theta}_{N})+\varepsilon/2\leq\hat{\mathcal{L}}_{N}(\theta^{*})+\varepsilon/2\leq\mathcal{L}(\theta^{*})+\varepsilon.$ (18)
Since the total number of configurations is $|\Theta|\leq\prod_{l=1}^{L}(2^{B})^{d_{l}(d_{l-1}+1)}$. The proof is complete by substituting $|\Theta|$.
$\Box$
Noise-Robust Hybrid Dynamics
@@
In this section, we confront the formidable optimization challenges stemming from the proliferation of constraints in large-scale networks in the aforementioned model. The abundance of constraints not only complicates the optimization process but also necessitates the determination of numerous penalty coefficients in the traditional penalty method. The task of figuring out these penalty coefficients becomes increasingly intricate as the number of constraints grows, leading to potential inefficiencies and suboptimal solutions.
@@
To surmount this hurdle, we use the Quantum Conditional Gradient Descent (QCGD) algorithm developed in [45], a cutting-edge hybrid classical-quantum framework tailored to address quadratic, linearly-constrained, binary optimization problems on quantum machines. Our workflow starts with data being fed into the neural network. The optimization problem is then formulated as a QCBO. This problem is subsequently linearized and transformed into a copositive programming problem. In the hybrid quantum-classical iterative process, during each iteration, the quantum conditional gradient method calculates the update direction by minimizing a linear approximation of the penalized proxy of the objective function. QCGD formulates the linear minimization problem as a QUBO sub-problem, which is suitable for Ising machines. CIM solves the QUBO problem, and the solution obtained is fed back into the classical computing step to calculate the direction of iteration. This iterative loop continues, with the QUBO matrix being updated at each step, until convergence towards an optimal solution is achieved.
In the following theorem, we prove that by sampling a certain number of solutions from the Ising machine at each iteration, the algorithm can still converge despite the presence of errors.
Theorem 3
Consider the Quadratic Constrained Binary Optimization (QCBO) problem solved using the Q-FW algorithm with a random quantum oracle. At step $t$, the error in the objective in a single run of the quantum oracle, has mean $\mu_{t}$ and variance $\sigma_{t}$, both of which have time-independent upper bounds. Then:
@@
    Q-FW algorithm converges to the optimal solution almost surely, i.e.,
@@
    The total time to solution for any fixed success probability satisfies:
    $\text{TTS}_{\text{QCBO}}=O\left(\tau\cdot T\cdot\log n\right),$ (21)
@@
  • $\tau$ is the time per quantum oracle call.
Proof: The subproblem in the $t$-th iteration is solved via $m_{t}=c\log n$ repetitions, selecting the best solution among samples $\{\mathbf{D}^{(i)}_{t}\}_{1\leq i\leq m_{t}}$, i.e., $\mathbf{D}_{t}=\mathbf{D}^{(i_{t})}_{t}$, where
@@
which is independent of iteration index $t$, assuming each element in $\mathbf{Q}^{(t)}_{\text{QUBO}}$ is constant. According to Proposition 1, we can get the order of convergence in expectation, and the almost sure convergence follows from the concentration probability bound that converges to $1$ when $T$ is infinity.
The bound on TTS can also be derived from the concentration result in Proposition 1. A more straightforward approach is based on the union bound, the probability that
@@
Therefore, With high probability $1-\frac{1}{\mathrm{poly}(n)}\geq p_{R}$, $\mathbf{D}_{t}=\mathbf{D}^{*}_{t}$ for all $t$, yielding the convergence to optimal solution. As per-iteration time of Q-FW is $m_{t}\tau$, we can obtain the second conclusion for $\text{TTS}_{\text{QCBO}}$. $\Box$
Quantum computers, such as those implementing adiabatic quantum computing (AQC) for solving QUBO problems, face challenges due to limited parameter precision [46]. In theory, QUBO parameters are real-valued, but practical hardware—whether quantum annealers like D-Wave’s or other accelerators—relies on finite precision representations, such as fixed-bit integers [46]. This limitation introduces perturbations in quantum systems, which distort the energy landscape and can shift the global optimum.
@@
Let the coefficients of the QUBO problem be truncated to $d=O(\log n)$ digits. If QCGD is used with these truncated coefficients, it will retain convergence guarantees and the number of iterations remains in the same order.
@@
The formal proof is presented in the supplementary. This result is appealing because it demonstrates that even with limited precision, the hybrid algorithm can still achieve convergence without a substantial increase in computational effort. This behavior can be interpreted as a form of implicit error correction. The introduced errors are effectively controlled and do not accumulate in a way that disrupts the convergence properties of the algorithm.
Experimental validation of quantum advantage
@@
To verify the advantages of the training on neural network, we undertake the experiment using coherent Ising machines for the classification of coat and sandal images from the Fashion MNIST dataset. The dataset consists of $60,000$ training images and $10,000$ test images, each representing a grayscale clothing item from $10$ classes, including T-shirts/tops, trousers, pullovers, dresses, coats, sandals, shirts, sneakers, bags, and ankle boots. Each image is $28\times 28$ pixels, with pixel values ranging from 00 to $255$. Additional details of the experimental results are deferred to the supplementary.
@@
Figure 3(a) illustrates the energy evolution during the optimization process using CIM to solve the neural network training directly. CIM identifies the optimal solution at the 88th iteration, approaching the optimal value in 0.185 ms, with the resulting model achieving a test accuracy of $94.95\%$.
@@
To evaluate the performance of our quantum-based approach for training quantized neural networks, we present a series of comparative analyses. Figure 3(b) compares the test accuracy and running time of our model and quantum approach against previous classical approaches, including straight through estimator (STE) [47] and BinaryConnect [48]. We provide a brief description of the two classical algorithms in the supplementary. Here the proposed quantum-based method achieves a test accuracy of $94.95\%$ with a significantly reduced running time of $0.185$ ms, demonstrating superior efficiency and effectiveness.
@@
Further, Figure 3(c) examines the memory demand and inference complexity, contrasting a neural network with continuous weights and sigmoid activation functions against our quantized approach, which approximates the activation function with a simple piecewise linear function. Our method reduces memory cost to 2.75 bytes and inference time to 0.0762 seconds while maintaining a test accuracy of $94.95\%$, highlighting its resource efficiency. These results collectively underscore the advantages of leveraging quantum computing for training quantized neural networks, offering improvements in speed, and resource efficiency.
Figure 3: (a) Energy evolution curve during the optimization of a quantized neural network using a coherent Ising machine (CIM); (b) Comparative analysis of test accuracy and running time for different training algorithms (STE [47], BinaryConnect [48]); (c) Performance comparison of memory demand, inference time, and accuracy before and after quantization and activation function approximation.
Table 2 summarizes the performance of post-training quantization (PTQ) algorithms under different activation functions and bit widths. As expected, reducing the bit precision generally leads to a degradation in accuracy.
Algorithm Activation FP Model 4-bit 3-bit 2-bit
RTN [49] ReLU 98.7% 95.4% 98.7% 50.0%
LeakyReLU 98.7% 98.7% 98.7% 75.4%
Sigmoid 98.7% 97.2% 95.4% 75.4%
AdaRound [9] ReLU 98.7% 98.7% 98.8% 50.0%
LeakyReLU 98.7% 98.7% 98.7% 98.7%
Sigmoid 98.7% 97.2% 95.4% 50.0%
DFQ [50] ReLU 98.7% 95.4% 98.7% 50.0%
LeakyReLU 98.7% 98.7% 98.7% 98.7%
Sigmoid 98.7% 98.7% 98.7% 50.0%
Table 2: Classfication accuracy (%) under different PTQ algorithms, activation functions, and bit widths.
While most methods maintain high accuracy at $4$-bit and $3$-bit precision, a significant drop is observed at $2$-bit, especially with ReLU activations. Specifically, our method achieves $94.95\%$ test accuracy with an effective parameter precision of just $1.1$ bits, substantially outperforming almost all listed $2$-bit baselines. This result highlights the potential of our framework under aggressive quantization.
Spline interpolation requires careful consideration of the number of intervals, as it directly influences qubit resource allocation and must be optimized for quantum hardware constraints. In our experimental setup, we partitioned the input range into four intervals, achieving an optimal trade-off between qubit efficiency and segmentation accuracy. The corresponding breakpoints, defined at $[-8,-c,0,c,8]$, yielded the following results:
@@
Figure 4 illustrates the convergence behavior of the Quantum Conditional Gradient Descent algorithm when applied to the neural network training utilizing a hybrid quantum-classical approach with CIM and Gurobi solvers. The algorithm achieves convergence to the optimal solution within 321 iterations, with the objective residual and constraint residual presented on a logarithmic scale. The CIM solver exhibits consistend convergence behavior, steadily driving both the objective and constraint residual below the threshold value. In contrast, the Gurobi solver maintains convergence but displays oscillatory patterns in the constraint residual between iterations 30-100. Notably, the hybrid strategy achieves convergence in just 321 iterations, demonstrating its computational efficiency by effectively combining the strengths of CIM and the classical approach. To further validate the robustness and reproducibility of these results, we performed five independent experiments achieving convergence to the optimal solution in 723, 338, 263, 432, and 427 iterations, respectively. These consistent outcomes across multiple trials confirms its reliability and reproducibility in practical applications.
Figure 5: (a) Circular heatmaps depict the evolution of the QUBO matrix during the QCGD algorithm for different significant digit precisions (1 to 6 digits) across selected iterations (300, 600, 900, 1200, 1500, 1800, 2100, 2400, 2700, and 3000). The color intensity represents matrix element values, with the color bar indicating the range from 0.0 (purple) to 1.0 (yellow). Each row corresponds to a specific precision level, illustrating the iterative trajectory and convergence behavior. (b) and (c): The two subplots show the smoothed absolute differences $|\mathrm{obj}_{k}-\mathrm{obj}_{10}|$ (left) and $|\mathrm{residual}_{k}-\mathrm{residual}_{10}|$ (right) relative to the 10-digit precision baseline across iterations for precision levels of 1 to 9 digits. The y-axis is on a log scale, with dashed lines representing each precision level.
Robustness of hybrid approach: impact of coefficient matrix rounding
To investigate the convergence behavior of the hybrid algorithm in neural network training under different matrix rounding precision levels, we systematically rounded the matrix coefficients to digit precisions ranging from 1 to 10 decimal places. The algorithm was then executed for each precision setting until convergence to the optimal solution was achieved. Specifically, matrix elements were rounded to 1 through 10 significant digits, and the corresponding number of iterations required to attain the optimal solution at each precision level was recorded as follows:
These results indicate the variability in convergence speed, with lower precision levels (e.g., 1-3 digits) requiring substantially more iterations (up to 43,343 for 3 digits), likely due to increased numerical instability and larger initial differences. In contrast, higher precision levels (e.g., 5-10 digits) converge more efficiently, with the minimum iterations observed at 2904 for 10 digits, suggesting that increased precision enhances the convergence stability.